OpenStudy (anonymous):

The two-way frequency table shows the results of a survey of college students. Find the probability that a student's first class of the day is a humanities class, given the student is male. Round to the nearest thousandth. First class of the day male female humanities 70 80 Science 50 80 other 60 70

OpenStudy (anonymous):

Probability of students first class of the day is a humanities given that the student is male. P( student male and first class of the day is humanities class) / P (student is male) P( student male and first class of the day is humanities class) = 70 / (70+80+50+80+60+70)= 0.17 P (student is male) = (70+50+60)/(70+80+50+80+60+70)= 0.439 Answer = 0.17 / 0.439 = 0.387